# Prompt-Guided Internal States for Hallucination Detection of Large Language Models

### v1

@@ 3.2 Effect of the Prompt on Structural Consistency
4 Methodology
The analyses in sections 3 indicate that the introduction of appropriate prompts can guide changes in the structure related to text truthfulness within the LLM’s internal states, and these changes will facilitate our use of LLMs’ internal states as features to perform the hallucination detection task. Based on this analysis, we propose a novel framework PRISM, i.e., PRompt-guided Internal States for hallucination detection of Large Language Models. This framework can be effectively integrated with existing hallucination detection methods to enhance their generalization across different domain data.
Our framework consists of two components. First, we manually construct a prompt template $P$ related to the hallucination detection task, and then we use the language model $L$ to generate multiple candidate prompt templates $P_{i}$ that are similar in meaning but different in structure. For each $P_{i}$, we combine it with text $a_{j}$ from the labeled dataset $A$ to obtain text $P_{i}(a_{j})$, which is then input into $L$ to obtain the corresponding feature vector $v_{ij}$. By using all feature vectors to calculate the structure salience index $R_{i}$ corresponding to each $P_{i}$ on $A$, we select the $P_{i}$ with the highest $R_{i}$ as our final prompt template $P_{s}$.
Next, we integrate the existing hallucination detection methods into our framework, training the hallucination detector $D$ on $A$ using all the feature vector $v_{j}$ and their label $b_{j}$ corresponding to the text $P_{s}(a_{j})$. For new text $T$ to be detected, we similarly combine it with the prompt template $P_{s}$ to obtain the text $P_{s}(T)$ and input it into the language model $L$ to obtain the corresponding feature vector $v$. Finally, we input $v$ into the hallucination detector $D$ to get its hallucination label $H$.
The above describes the complete process of our framework. Next, we will provide a more detailed explanation of some key steps in this framework.
@@ 4.1 Prompt Generation
We need to generate a large number of suitable candidate prompt templates for the hallucination detection task, and we can use LLMs to accomplish this process. First, we manually construct a simple prompt, such as Prompt 1 in Section 3. Then, we use this prompt to construct the following prompt:
@@
Finally, we input the above prompt into the LLM to obtain a large number of prompt templates similar to Prompt 1. For example, we used Prompt 2 to query GPT-4o, resulting in 10 prompt templates that have similar meanings but different formats, with the relevant content presented in Appendix C.
@@
4.2 Prompt Selection
It is worth noting that this process relies on the selection of feature vectors from the internal states of the LLM. We selected the context embedding corresponding to the last token in the final layer of LLaMA2-Chat-7B as the feature vector and calculated the corresponding variance ratio for each of the 10 prompt templates from Appendix C. We computed the variance ratio for each prompt template across the six sub-datasets of the True-False dataset, averaged the results, and ranked them, as shown in Table 2.
From Table 2, we can see that all the prompt templates increased the variance proportion along the truthfulness direction in the True-False dataset. Among them, the 10th prompt template achieved the highest variance ratio and it is as follows:
Prompt 3:
Does the statement ’[statement]’ accurately
reflect the truth?
Therefore, we will select this prompt template for our framework in the experiments presented in the next section.
@@ 5 Experiment
5.1 Dataset and Metric
We evaluate our proposed framework PRISM and other baselines on the True-False and LogicStruct datasets which are detailed in section 2. Since the ratio of the two classes of labels in these datasets is close to 1:1, classification accuracy is an appropriate metric.
@@ 5.2 Baselines
We choose the following hallucination detection methods as baselines:
  • LN-PP Manakul et al. (2023) detects hallucinations based on the average probabilities of tokens generated by the LLM. We first determine a threshold, and then use this threshold for label classification.
  • SAPLMA Azaria and Mitchell (2023) is an effective supervised learning method. It uses activation values within the LLM as feature vectors to train a fully connected network as the hallucination detector. Selections of activation values from different layers can all achieve good classification performance.
  • MIND Su et al. (2024) is an unsupervised learning framework. It automatically generates labeled datasets from high-quality Wikipedia articles for training the hallucination detector. Other settings are consistent with the SAPLMA method.
  • MMMarks and Tegmark (2023) uses a simple mass-mean probe to distinguish between true and false statements. This method selects specific activation values as feature vectors, then uses Equation (1) to calculate the truthfulness direction on the labeled dataset. For the text to be detected, it first projects the feature vector onto the truthfulness direction by taking the dot product, then applies the sigmoid function, and finally rounds the result to obtain the label.
@@ 5.3 Implementation Details
We conducted our experiments using LLaMA2-7B-Chat and LLaMA2-13B-Chat Touvron et al. (2023). The activation values were selected from the context embeddings corresponding to the last token in the last layer of the LLMs. In the SAPLMA and MIND methods, the dimensions of the detectors for each layer were 256, 128, 64, and 2, with a dropout rate of 20% applied in the first layer. The activation function was ReLU, and the optimizer was Adam.
Our experiments focused on the generalization performance of different methods across different domains. For the SAPLMA method, we trained the detector separately on each subset of the True-False dataset and tested it on all other subsets. Thus, for a single subset, we obtained multiple test results from detectors trained on other subsets and averaged these values to obtain the final accuracy for this subset. In the LogicStruct dataset, we concentrated on training with affirmative statements and testing on other grammatical structures, as previous studies have indicated that achieving good generalization results in this scenario is challenging Marks and Tegmark (2023); Bürger et al. (2024). Both the MM and LN-PP methods employed the same evaluation approach as SAPLMA. The MIND method was trained on its automatically generated dataset and tested on each subset of the True-False and LogicStruct datasets. The settings of the methods under PRISM framework remain consistent with those of the original methods.
@@ 5.4 Experimental Results
We can see from Table 3 and 4 that our framework significantly outperforms other baselines in terms of average accuracy on the True-False dataset. On almost every sub-dataset of different topics, our framework substantially improves the generalization performance of the original methods. Additionally, there is a significant difference in the performance of the original MM and SAPLMA methods, which may be due to the simpler structure of the detector used in the MM method compared to that in the SAPLMA method. However, after integrating both into our framework, these two methods achieved similar performance. This indicates that the introduction of prompts made the structure related to text truthfulness more salient in the dataset, making it easier for detectors to capture. On the Facts dataset, the SAPLMA method and our framework achieve nearly identical accuracy. As shown in Figure 1, before the introduction of prompts, the true and false statements on the Facts dataset already exhibit clear separability, making it difficult to further enhance this structure.
Table 5 presents the experimental results on the LogicStruct dataset. We can observe that, except for our framework, other baselines struggle to generalize the training results on affirmative statements to other grammatical structures. According to the observations made by Bürger et al. (2024), this may be because the structure related to text truthfulness in affirmative statements is different from the corresponding structure in other grammatical structures. In contrast, our framework achieves significant classification performance across these grammatical structures, indicating that the introduction of prompts indeed makes the structure related to text truthfulness more consistent across different datasets. Additionally, we observe that the average results on the 13B model outperform those on the 7B model, suggesting that as the model’s reasoning capabilities improve, our framework is likely to perform better as well.
@@ 5.5.1 Impact of Prompt Selection
We conducted empirical experiments using the PRISM-SAPLMA method on all 10 prompt templates in Appendix C, with the results presented in Table 6. We can observe that all the prompt templates significantly improved the average accuracy compared to the SAPLMA method in Table 3, indicating that the improvement in the generalization performance of the hallucination detector due to the prompts is general and stable.
Additionally, by combining the ranking information from Table 2 and Table 6, we can see that the optimal prompt template selected based on the variance ratio also achieved the best results in the actual experiments. Furthermore, the 7th prompt template performs significantly worse in the actual experiments compared to others, while we ranked it 9th based on the variance ratio. These results indicate that the variance ratio can help us select prompt templates that achieve better generalization results in actual experiments while avoiding poorly performing ones.
@@ 5.5.2 Impact of Layer Selection
In addition to selecting context embeddings corresponding to the last token in the last layer of LLMs as feature vectors, we also extracted embeddings corresponding to the last token from the middle layer (16th) of LLaMA2-7B-Chat. We conducted experiments on the True-False dataset with our framework and the corresponding original methods, keeping all other settings the same, and the results are as follows:
Baselines Average
MM 0.5166
SAPLMA 0.5766
PRISM-MM 0.7924
PRISM-SAPLMA 0.7938
Compared to Table 3, the performance of the original methods has worsened, while our framework has achieved better results. This indicates that our framework exhibits good stability with respect to different layer selections.
Table 6: The experimental results of all 10 prompt templates in Appendix C on the True-False dataset using LLaMA2-7B-Chat.
Prompt Animals Cities Comp. Elements Facts Invent. Average Ranking
I 0.6889 0.8823 0.8302 0.7066 0.7475 0.7721 0.7713 2
II 0.6395 0.8915 0.8138 0.6318 0.7419 0.7183 0.7395 7
III 0.6460 0.9021 0.8087 0.6394 0.6956 0.7016 0.7322 8
IV 0.6887 0.8959 0.8058 0.6869 0.7436 0.7797 0.7668 3
V 0.7004 0.8506 0.8005 0.6639 0.7550 0.7123 0.7471 6
VI 0.6895 0.9115 0.7903 0.6699 0.7396 0.7587 0.7599 5
VII 0.6540 0.7764 0.7053 0.6200 0.7406 0.6532 0.6916 10
VIII 0.6829 0.9239 0.8205 0.6809 0.7318 0.7347 0.7624 4
IX 0.6310 0.8594 0.7979 0.6303 0.7031 0.7011 0.7205 9
X 0.7133 0.8930 0.8297 0.6736 0.7501 0.7801 0.7733 1
5.5.3 Impact of the Internal States
After introducing the prompt, is it possible to assess the truthfulness of the text directly based on the LLM’s response, without relying on its internal states? To investigate this question, we directly extract the token probabilities for " Yes" and " No" from LLaMA2-7B-Chat when generating the next token and compute their ratio $p[3869]/p[1939]$, where $p$ represents the vocabulary probabilities for generating the next token. We use whether this ratio is greater than or equal to 1 to determine the label of the original text. The average accuracy on the True-False dataset is as follows:
Baselines Average
@@ 5.5.3 Impact of the Internal States
PRISM-MM 0.7524
PRISM-SAPLMA 0.7735
We can observe a significant accuracy difference between directly using the responses generated by the LLM and training the detector based on the LLM’s internal states. This indicates that the LLM’s internal states contain more information related to text truthfulness, which is not observable from the LLM’s responses but still plays a crucial role in hallucination detection.
@@ 6 Conclusions
In this paper, we first investigate the effect of specific prompts on the structure related to text truthfulness within the internal states of LLMs. We find that suitable prompts can make this structure more salient and consistent across datasets from different domains, facilitating hallucination detectors to learn this structure and improve generalization. Based on this insight, we introduce a novel framework, prompt-based internal states for hallucination detection of large language models, namely PRISM. This framework initially generates suitable prompt templates for hallucination detection task through a simple and effective approach, then integrates with existing hallucination detection methods to achieve detectors with improved generalization. We conducted experiments across various baselines on datasets from different domains, and the results demonstrate that our framework significantly enhances the generalization performance of existing methods and remains stable with different prompt and feature selections. We hope our framework can inspire future research in LLMs hallucination detection.
Limitations
We acknowledge that our research still has certain limitations. When detecting hallucinations in content generated by LLMs, we rely on the model’s internal states, which can be challenging to access in some cases and may require the use of a proxy model. Additionally, our approach does not leverage other information produced by the LLM during text generation, such as probability data and the generated text itself. In future research, we will explore fully utilizing various types of information produced by LLMs during text generation to achieve more effective hallucination detection.
Ethics Statement
The development and application of the PRISM framework for hallucination detection in large language models (LLMs) are guided by the principle of assessing the reliability and credibility of AI-generated content. PRISM aims to reduce the risk of users being misled by AI-generated outputs, thereby contributing to safer and more trustworthy AI applications. Additionally, the datasets used in this project are publicly available and do not involve personal or sensitive data, ensuring privacy and security. We need to recognize that our detector classifies based on the internal states of the model, and thus may inherit biases from the language model on certain issues, potentially leading to incorrect judgments by the detector. Therefore, we encourage further research and collaboration to develop ethical, fair, and trustworthy AI systems.
@@ Appendix A Dataset Details
In both logical conjunctions and logical disjunctions, Statement 1 and Statement 2 are derived from affirmative statements and sampled in a way that ensures label balance across the overall dataset. Additional information is presented in Table 8.
@@ Appendix B Analysis of Prompt 3
Consistent with the analysis of Prompt 1 in Section 3, we computed the ratio of the variance along the truthfulness direction to the total variance for the six sub-datasets of the True-False dataset. The results before and after the introduction of Prompt 3 are presented in Figure 3. We also computed the cosine similarity of the truthfulness direction between these datasets after the introduction of Prompt 3, as shown in Table 9.
Figure 3: Comparison of variance ratios before and after the introduction of Prompt 3 on the True-False dataset.
We can observe that Prompt 3 significantly enhances the salience and consistency of the structure related to text truthfulness in the embeddings. This is consistent with the analysis of Prompt 1 in Section 3, indicating that using prompts to guide changes in this structure within the LLM’s internal states is generally effective.
Table 9: Cosine similarity between truthfulness directions across different topics in the True-False dataset after Prompt 3 addition.

### v2

@@ 4.1 Prompt Generation
To obtain suitable candidate prompt templates for the hallucination detection task, we can use LLMs to assist in this process. First, we manually construct a simple prompt, such as Prompt 1 in Section 3. Next, we use this prompt to construct the following one:
@@
Finally, we input the above prompt into LLMs to obtain a large number of prompt templates similar to Prompt 1. For example, we used Prompt 2 to query GPT-4o, resulting in 10 prompt templates that have similar meanings but different formats. The relevant content is provided in Appendix C.
4.2 Prompt Selection
It is worth noting that this process relies on the selection of feature vectors from the internal states of LLMs. We selected the contextualized embedding corresponding to the last token in the final layer of LLaMA2-Chat-7B as the feature vector and calculated the corresponding variance ratio for each of the 10 prompt templates in Appendix C. We computed the variance ratios for each prompt template across the six sub-datasets of the True-False dataset, averaged the results, and ranked them, as shown in Table 2.
According to Table 2, the 10th prompt template achieved the highest variance ratio and we present its form below:
Prompt 3:
Does the statement ’[statement]’ accurately
reflect the truth?
We selected this prompt template for our framework in the experiments presented in Section 5. We also conducted the same analysis on this prompt as we did in Section 3 and presented the results in Appendix B.
@@ 5 Experiment
5.1 Datasets and Metric
We evaluate our PRISM and other baselines on the True-False and LogicStruct datasets which are detailed in section 2. Since the ratio of the two classes of labels in these datasets is close to 1:1, classification accuracy is an appropriate metric.
@@ 5.2 Baselines
We choose the following methods as baselines:
  • LN-PP Manakul et al. (2023) detects hallucinations based on the average probabilities of tokens generated by LLMs. We first determine a threshold, and then use it for label classification.
  • SAPLMA Azaria and Mitchell (2023) is an effective supervised learning method. It uses activation values in LLMs as feature vectors to train a Multilayer Perceptron (MLP) network as the hallucination detector.
  • MIND Su et al. (2024) is an unsupervised learning framework. It automatically generates labeled datasets from Wikipedia articles for training the hallucination detector. Other settings are consistent with the SAPLMA method.
  • MMMarks and Tegmark (2023) uses a simple mass-mean probe to distinguish between true and false statements. This method selects specific activation values as feature vectors and uses Equation (1) to calculate the truthfulness direction on the labeled dataset. For the text to be detected, it first projects the feature vector onto the truthfulness direction by taking the dot product, then applies the sigmoid function, and finally rounds the result to obtain the label.
@@ 5.3 Implementation Details
We conducted our experiments using LLaMA2-7B-Chat and LLaMA2-13B-Chat Touvron et al. (2023b). The activation values were selected from the contextualized embeddings corresponding to the last token in the final layer of the LLMs. In the SAPLMA and MIND methods, the dimensions of the detectors for each layer were 256, 128, 64, and 2, with a dropout rate of 20% applied in the first layer. The activation function was ReLU, and the optimizer was Adam. These settings are the same as in the original paper Su et al. (2024).
Our experiments focused on the generalization performance of different methods across different domains. For the SAPLMA method, we trained the detector separately on each sub-dataset of the True-False dataset and tested it on all other sub-datasets. Thus, for a single sub-dataset, we obtained multiple test results from detectors trained on other sub-datasets, and we averaged these results to obtain the final classification accuracy for this sub-dataset. In the LogicStruct dataset, we concentrated on training with affirmative statements and testing on other grammatical structures, as previous studies have indicated that achieving good generalization results in this scenario is challenging Marks and Tegmark (2023); Bürger et al. (2024). Both the MM and LN-PP methods employed the same evaluation approach as the SAPLMA. The MIND method was trained on its automatically generated dataset and tested on each sub-dataset of the True-False and LogicStruct datasets. The settings of the methods under PRISM framework remain consistent with those of the original methods.
@@
Table 6: The experimental results of all 10 prompt templates on the True-False dataset using LLaMA2-7B-Chat.
Prompt I II III IV V VI VII VIII IX X Without
Acc. 0.7217 0.7678 0.7297 0.7398 0.6997 0.7749 0.6931 0.7162 0.7003 0.7524 0.5572
Ranking 6 2 5 4 9 1 10 7 8 3 11
@@ 5.4 Experimental Results
We can see from Table 3 and 4 that our framework significantly outperforms other baselines in terms of the classification accuracy on the True-False dataset. On almost every sub-dataset of different topics (5 out of 6 topics), our framework substantially improves the generalization performance of the original methods. Additionally, there is a significant difference in the performance of the original MM and SAPLMA methods, which may be due to the simpler structure of the detector used in the MM method compared to that in the SAPLMA method. However, after integrating both into our framework, these two methods achieve similar performance. According to Section 3.1, this may be because the introduction of the prompt makes the structure related to text truthfulness more salient in the LLMs’ internal states, making it easier for detectors to capture this structure. On the Facts sub-dataset, the original SAPLMA method and our framework achieve nearly identical accuracy. As shown in Figure 1, before the introduction of prompts, the true and false statements in the Facts sub-dataset already exhibit clear separability, so it is difficult to further enhance this structure by appropriate prompts.
Table 5 presents the experimental results on the LogicStruct dataset. We can observe that, except for our framework, other baselines struggle to generalize the training results on affirmative statements to other grammatical structures. According to the observations made by Bürger et al. (2024), this is because the structure related to text truthfulness in affirmative statements is different from the corresponding structure in other grammatical structures. However, our framework achieves significant generalization performance across different grammatical structures. This indicates that the introduction of the prompt indeed makes the structure related to text truthfulness more consistent across different datasets, which aligns with the observation in Section 3.2.
@@ 5.5.1 Impact of Prompt Selection
In this section, we focus on the impact of prompt selection. We conducted experiments using the PRISM-MM method on all 10 prompt templates presented in Appendix C, with the results presented in Table 6. We can observe that all the prompt templates significantly improved the average accuracy compared to the result of the original MM method. This indicates that the impact of prompts on improving the generalization performance of hallucination detectors is general and stable.
In addition, by combining the ranking information from Table 2 and Table 6, we can see that the top 4 prompt templates selected based on the variance ratios all perform within the top 5 in actual experiments, while the two worst-performing prompt templates correspond to the two lowest variance ratios. These results indicate that the variance ratio can help us select prompt templates that achieve better generalization results in actual experiments while avoiding poorly performing ones.
@@ 5.5.2 Impact of Layer Selection
In addition to selecting the contextualized embeddings corresponding to the last token in the final layer of LLMs as feature vectors, we also extracted the embeddings corresponding to the last token in the middle layer (16th) of LLaMA2-7B-Chat to evaluate our framework. We conducted the same experiments on the True-False dataset with our framework and the corresponding original methods, as shown in the last four rows of Table 7.
Table 7: The experimental results on the True-False datase using feature vectors from the middle layer of LLaMA2-7B-Chat and the next token probabilities from LLaMA2-7B-Chat.
Baselines Average Accuracy
@@
MM(middle) 0.5166
SAPLMA(middle) 0.5766
PRISM-MM(middle) 0.7924
PRISM-SAPLMA(middle) 0.7938
Compared to the results in Table 3, the performance of the original methods has worsened, while our framework has achieved better results. This indicates that our framework exhibits good stability with respect to different layer selections.
5.5.3 Impact of Internal States
After introducing the prompt, is it possible to assess the truthfulness of the text directly based on LLMs’ responses without relying on their internal states? To investigate this question, after inputting the prompt into LLaMA2-7B-Chat, we directly extracted the token probabilities for " Yes" and " No" when the LLM was going to generate the next token and calculated their ratio $p[3869]/p[1939]$, where $p$ represents the vocabulary probabilities for generating the next token. We used whether the ratio is greater than 1 to determine the label of the original text. The average accuracy on the True-False dataset is shown in the first row of Table 7.
We can observe a significant accuracy difference between directly using the responses generated by the LLM and training detectors based on the prompt-guided internal states. This indicates that the LLM’s internal states contain more information related to text truthfulness, which is not observable from the LLM’s responses but still plays a crucial role in hallucination detection.
@@ 6 Conclusions
In this paper, we first investigated the effect of specific prompts on the structure related to text truthfulness in LLMs’ internal states. We found that suitable prompts can make this structure more salient and consistent across datasets from different domains, facilitating hallucination detectors in learning this structure and improving their generalization performance. Based on this insight, we introduced a novel framework, prompt-guided internal states for hallucination detection of large language models, namely PRISM. This framework integrates the prompt-guided internal states with existing hallucination detection methods to obtain more advanced detectors. Finally, we conducted experiments across various baselines on datasets from different domains, and the results demonstrated that our framework can significantly enhance the generalization performance of existing hallucination detection methods.
7 Limitations
We acknowledge that our research still has some limitations. When detecting hallucinations in the content generated by LLMs, we heavily rely on their internal states, which may be challenging to access in some cases, requiring us to use a proxy model. Additionally, our framework does not leverage other information produced by LLMs during text generation, such as probability information and the generated text itself. In future research, we will explore fully utilizing various types of information produced by LLMs during text generation to achieve more effective hallucination detection.
8 Ethics Statement
The development and application of the PRISM framework for Large Language Models hallucination detection are guided by the principle of assessing the reliability and credibility of AI-generated content. PRISM aims to reduce the risk of users being misled by AI-generated content, thereby contributing to safer and more trustworthy AI applications. Additionally, the datasets used in this project are publicly available and do not involve personal or sensitive data, ensuring privacy and security. We need to recognize that our framework classifies text based on the model’s internal states, and thus may inherit biases from the language model on certain issues, potentially leading to incorrect judgments. Therefore, we encourage further research and collaboration to develop ethical, fair, and trustworthy AI systems.
@@ Appendix A Dataset Details
In both Logical conjunctions and Logical disjunctions, statement 1 and statement 2 are derived from Affirmative statements and sampled in a way that ensures label balance across each dataset. Additional information is presented in Table 9.
@@ Appendix B Analysis of Prompt 3
Consistent with the analysis of Prompt 1 in Section 3, we calculated the variance ratios of Prompt 3 on each sub-dataset of the True-False dataset. The results before and after the introduction of Prompt 3 are presented in Figure 3. We also calculated the cosine similarity between the truthfulness directions across these sub-datasets after the introduction of Prompt 3, as shown in Table 10.
We can observe that Prompt 3 significantly enhances the salience and consistency of the structure related to text truthfulness in the embeddings. This is consistent with the conclusion in Section 3, indicating that using prompts to guide changes in LLMs’ internal states is generally effective.
Table 10: The cosine similarity between the truthfulness directions across different topics in the True-False dataset after Prompt 3 addition.
@@
Figure 3: Comparison of the variance ratios before and after the introduction of Prompt 3 on each sub-dataset of the True-False dataset.